Solve the system by the elimination method:
5X-Y=5 AND 3X+5Y=-2

Answer by mananth(16946)

5 x -1 y = 5 .............1
Total value
3 x + 5 y = -2 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 1
25 x -5 y = 25
3 x + 5 y = -2
Add the two equations
28 x = 23
/ 28
x = 23/28
plug value of x in (1)
5 x -1 y = 5
4 3/28 -1 y = 5
-1 y = 5 -4 3/28
-1 y = 25/28
y = - 25/28
x= 23/28
y= - 25/28